A processing station was purchased by a company for $10,000. After 4 years it is estimated that the value of the station will be $4,000. If the value in dollors V and the time the station has been in use t are related by a linear equation, find the equation that relates v AND T
Answer by stanbon(75887)

A processing station was purchased by a company for $10,000. After 4 years it is estimated that the value of the station will be $4,000. If the value in dollors V and the time the station has been in use t are related by a linear equation, find the equation that relates v AND t
----------------------------------
You have two points relating v and t: (0,10000) and (4,4000)
slope = (4000-10000)/(4-0) = -6000/4 = -1500
intercept: 10000
--------------------
EQUATION:
v(t) = -1500t + 10,000
=========================
Cheers,
Stan H.